3. Calculate the moment of force F = 5 N, for the case when d1 = 20 cm and d2 = 80 cm. How many times does the rotational effect of the force F increase in the second case compared to the first?


1
Expert's answer
2021-02-22T10:26:52-0500

(a)

"M_1=Fd_1=5\ N\cdot0.2\ m=1.0\ N\cdot m,""M_2=Fd_2=5\ N\cdot0.8\ m=4.0\ N\cdot m."

(b)

"\dfrac{M_2}{M_1}=\dfrac{4.0\ N\cdot m}{1.0\ N\cdot m}=4."


The rotational effect of the force F in the second case 4 times larger compared to the first.